### v1

# A Novel Deep Hybrid Framework with Ensemble-Based Feature Optimization for Robust Real-Time Human Activity Recognition

Abstract
Human Activity Recognition (HAR) plays a pivotal role in various applications, including smart surveillance, healthcare, assistive technologies, sports analytics, etc. However, HAR systems still face critical challenges, including high computational costs, redundant features, and limited scalability in real-time scenarios. An optimized hybrid deep learning framework is introduced that integrates a customized InceptionV3, an LSTM architecture, and a novel ensemble-based feature selection strategy. The proposed framework first extracts spatial descriptors using the customized InceptionV3 model, which captures multilevel contextual patterns, region homogeneity, and fine-grained localization cues. The temporal dependencies across frames are then modeled using LSTMs to effectively encode motion dynamics. Finally, an ensemble-based genetic algorithm with Adaptive Dynamic Fitness Sharing and Attention (ADFSA) is employed to select a compact and optimized feature set by dynamically balancing objectives such as accuracy, redundancy, uniqueness, and complexity reduction. Consequently, the selected feature subsets, which are both diverse and discriminative, enable various lightweight machine learning classifiers to achieve accurate and robust HAR in heterogeneous environments. Experimental results on the robust UCF-YouTube dataset, which presents challenges such as occlusion, cluttered backgrounds, motion dynamics, and poor illumination, demonstrate good performance. The proposed approach achieves 99.65% recognition accuracy, reduces features to as few as 7, and enhances inference time. The lightweight and scalable nature of the HAR system supports real-time deployment on edge devices such as Raspberry Pi, enabling practical applications in intelligent, resource-aware environments, including public safety, assistive technology, and autonomous monitoring systems.
Keywords: Human Activity Recognition, Hybrid Deep Learning, InceptionV3, LSTM Networks, Ensemble Feature Selection, Adaptive Dynamic Fitness Sharing, Real-Time Processing, Edge Computing
1 Introduction
Human activity recognition (HAR) plays a pivotal role in the development of smart cities, particularly in the context of video-based surveillance. Accurate identification of human behavior can help law enforcement and public safety agencies detect irregular activities, such as accidents, property damage, or illicit behaviors such as fighting and theft. Early research in HAR focused primarily on controlled environments with a single actor, but recent studies have shifted toward more realistic and challenging scenarios. These include uncontrolled video datasets with complex backgrounds, varying illumination, camera movement, and occlusion, making accurate recognition considerably more difficult [1]. Improving pedestrian detection and tracking in such dynamic surveillance situations remains a significant challenge.
Although Convolutional Neural Networks (CNNs) have demonstrated remarkable performance in image classification and object detection, they are inherently limited to processing individual frames. They cannot directly capture temporal dynamics in video sequences [2]. To address this, two-stream CNNs were proposed, in which appearance and motion information (e.g., optical flow) are processed separately. Although they achieve good recognition performance, they incur high computational costs, limiting their suitability for large-scale or real-time applications [3, 4]. More recently, 3D CNNs have been introduced to directly extract spatiotemporal information from raw video streams [5, 6]. To enhance temporal modeling, recurrent architectures such as RNNs and LSTMs have also been incorporated [7, 8]. These approaches suffer from high computational cost, especially when processing long video sequences, which limits their applicability for large-scale and real-time scenarios.
To address these limitations, researchers have explored feature fusion and selection strategies for HAR. Examples include combining deep spatial and motion representations using spatial–temporal attention mechanisms [9] and combining handcrafted multiview features with feature selection based on entropy-variance [10]. Although these methods report promising performance, they often rely on hand-crafted components and produce high-dimensional feature representations, leading to increased computational complexity and limited scalability. Efficient selection of discriminative features is therefore essential to improve inference speed without compromising accuracy.
In this work, we propose a novel hybrid HAR framework that integrates a customized InceptionV3 CNN for spatial feature extraction with an LSTM network for temporal modeling. The CNN captures discriminative spatial patterns from individual frames, while the LSTM models temporal dependencies across sequences. Designed for real-time video processing, the framework handles challenges such as occlusion, cluttered backgrounds, motion blur, and multiview variations, as exemplified by the UCF YouTube action dataset [11]. Both models are end-to-end trainable and optimized via transfer learning and fine-tuning to ensure efficient inference. Furthermore, to reduce feature redundancy and computational overhead, a genetic algorithm (GA)–based feature selection strategy is applied, selecting only the most discriminative spatiotemporal features for final classification. By eliminating the need for costly preprocessing steps such as optical flow or skeleton extraction, the proposed approach achieves high accuracy while maintaining practical real-time applicability.
The following are the main contributions of this work:
  • This study proposes a hybrid DL framework that synergistically integrates spatial customized InceptionV3 and temporal (LSTM) modeling for HAR from video data. The customized InceptionV3 integrates region-based (average pooling) and boundary-aware (max pooling) operations after each convolutional block to enhance region homogeneity, suppress noise, and capture discriminative local features, while improving robustness through down-sampling. Moreover, the enhanced InceptionV3 extracts fine-grained localization, contextual cues, and human positioning within frames, which a customized LSTM sequentially processes to model inter-frame temporal dependencies and motion dynamics with high fidelity.
  • A novel ensemble-based GA-driven feature selection strategy, ADFSA, is employed to achieve compact and discriminative feature subsets. The ADFSA mechanism dynamically reweights multiple objectives: classification accuracy, redundancy minimization, uniqueness enhancement, and complexity reduction through attention-guided adaptation, thereby accelerating convergence and mitigating local stagnation during the evolutionary process.
  • The optimally selected and diverse feature subsets include both spatial and temporal, are highly compact, diverse, and discriminative, enabling lightweight ML classifiers to achieve robust HAR performance across heterogeneous conditions with minimal resource overhead.
  • The proposed framework outperforms state-of-the-art approaches on the UCF–YouTube dataset with 99.65% recognition accuracy, while reducing the feature space from 128 to 7. This compact representation enables real-time inference on edge devices such as a Raspberry Pi. .
The rest of the paper is organized as follows. Section II provides related work. Section III presents the methodology and proposes the ADFSA scheme. Section IV is related to the experimental setup. Results and discussion are elaborated in section V. Section VI concludes the paper.
2 Related Work
Recent studies show a growing interest in leveraging advanced deep learning architectures for human action recognition, including CNN-based models, transfer learning approaches, and, more recently, Transformer-based sequence learning. For example, the BT-LSTM model proposed in [12] achieved an accuracy of 85.3% on the UCF11 dataset. Considerable progress has since been made with various deep learning approaches that attempt to jointly capture spatial and temporal information. Early methods such as the CNN–BiLSTM network [13] and the deep autoencoder combined with CNN [14] reported recognition accuracies of 92.84% and 96.2%, respectively. However, these end-to-end architectures require large-scale training datasets and entail high computational costs.
To improve spatial representation, subsequent studies explored handcrafted or hybrid descriptors. For instance, the key-frame driven KFDI approach [15] achieved only 79.4% due to limited spatial discrimination. Although more advanced techniques, such as the dilated CNN–BiLSTM with residual blocks [16] and the local–global feature fusion method with QSVM [17], achieved 89.01% and 82.6% accuracy, respectively, their generalization capability remained limited when applied to more complex and unconstrained activities.
More recently, fully 3D networks and transformer-based encoders have been introduced to capture long-term temporal dependencies. The 3D‐CNN in [18] achieved 85.20%, and the ConvNeXt–TCN model in [19] reached 97.73%, demonstrating the importance of temporal aggregation. Similarly, hybrid deep models such as VGG–BiGRU [20] and ViT‐ReT [21] achieved 93.38% and 92.4%, respectively. Despite these advances, most of these approaches employ complex, parameter-heavy architectures and retain high-dimensional feature representations, which leads to increased inference time and reduced efficiency.
Although the bidirectional LSTM model presented in [22] achieved a high recognition accuracy of 99.2%, it still relies on full feature vectors without any dimensionality reduction, limiting its suitability for real-time or edge deployment. In contrast, the proposed method attains a higher accuracy of 99.58% while substantially reducing the feature dimensionality. By extracting spatial features using a lightweight CNN and subsequently applying efficient machine learning classifiers to a compact set of selected features, the proposed framework provides a robust and computationally efficient solution that outperforms existing methods in both accuracy and inference efficiency. A summary of reported previous work is presented in Table 1.
Table 1: Summary of Previous Works on Human Activity Recognition
Ref. Dataset(s) Results / Method Key Shortcomings
[12] UCF11, KTH Block-term tensor RNN; 94% (UCF11) Requires tensor decomposition + full RNN; high computation
[13] UCF11, KTH CNN features + BiLSTM; 94.9% No feature reduction → high-dimensional inputs → slower inference
[14] UCF11, surveillance Deep autoencoder + CNN; 93.6% Lacks explicit temporal modeling; not optimized for real-time
[15] UCF11 Dynamic Image + CNN; 95.7% Ignores temporal sequences → limited spatio-temporal modeling
[16] UCF11, HMDB51, UCF50 Dilated CNN + Attention-LSTM; 96.5% (UCF11) Heavy attention block; poor generalization (68.4% HMDB51)
[17] EUCF-11 / RUCF-11 Spiking CNN for event data; 92.5% / 95.7% Designed for neuromorphic data, not RGB video; no feature reduction
[18] UCF101 3D-CNN; 94.1% High memory and inference time due to 3D convolutions
[19] UCF11, UCF50, UCF101, JHMDB ConvNeXt + TCN; 97.73%, 98.81%, 98.46%, 83.38% Very accurate, but deep/heavy model; no feature compression
[20] UCF Sports, JHMDB, UCF101 CNN + BiGRU + FS; 95.2%, 78.7%, 93.4% Features still high-dimensional; no real-time deployment
[21] UCF101, HMDB51 ViT-ReT; 97.4%, 82.6% Transformer complexity; high inference cost
[22] UCF11, JHMDB TL + BiLSTM; 96.7%, 79.5% Uses full feature vectors; no dimensionality reduction
To overcome these shortcomings, this study proposes an ensemble ADFSA approach. The method integrates multiple optimization objectives, including accuracy, redundancy, uniqueness, and complexity, that dynamically adjust their relative importance throughout the evolutionary process. Additionally, by leveraging a two-stage deep feature extraction pipeline (customized InceptionV3 + LSTM), the method effectively captures both spatial and temporal representations. This hybrid design enables the proposed approach to achieve better recognition performance, faster inference, and reduced model complexity, making it well-suited for real-time HAR applications.
3 Methodology
This section introduces the proposed novel hybrid deep feature extraction and selection methodology for HAR. In video-based HAR systems, human activities are mainly recognized from continuous video streams acquired through surveillance cameras or other visual sensors. The overall proposed framework for hybrid deep feature extraction and selection is shown in Figure 1, and the detailed explanation and presentation are provided in subsequent sections.
Figure 1: Overall Block Diagram of Proposed Methodology.
3.1 Frames Extraction from Video Clip
As deep CNN operate on image data, it is essential to extract individual frames from the video stream. To achieve this, a windowing mechanism is employed, which determines the interval at which frames are sampled and stored. This interval, referred to as the skip window size or sampling interval, is calculated using the following formula:
$windowsize=\text{max}\left(\frac{\text{video\_frames\_count}}{\text{Sequence\_Length}}\right).$ (1)
Figure 2: Video to Frames Extraction.
Initially, video data should undergo a frame extraction process, where individual frames are captured at a predetermined rate and resolution, as illustrated in Figure 2, in order to obtain a collection of images with uniform characteristics throughout the dataset.
3.2 Pre-processing
In this study, image preprocessing is applied to align the input dimensions of video frames with the expected input format of the modified deep convolutional architecture. Each frame is resized to $224\times 224$ pixels with 3 color channels (RGB), ensuring compatibility with the customized Inception-based feature extractor. Alongside resizing, pixel intensities are normalized to a fixed scale, typically in the range $[0,1]$, or standardized using the appropriate preprocessing function. This step harmonizes the input distribution with that of the network’s training conditions, promoting stable convergence and effective spatial feature extraction from each frame in the sequence.
3.3 Data Augmentation
To enhance the diversity and generalization of the model, data augmentation techniques are applied during the pre-processing phase. These augmentations simulate real-world variations and prevent overfitting by introducing randomness into the training data. Common transformations include random horizontal flipping, slight rotations, zooming, and brightness adjustments, all applied within controlled bounds to preserve the semantic content of human activities. These augmentations are executed on a per-frame basis before the frames are fed into the feature extractor. By increasing variability in spatial patterns without altering temporal coherence, data augmentation enhances the robustness and reliability of temporal feature learning in downstream LSTM-based modeling.
3.4 Deep Hybrid Convolutional Neural Network
This paper proposes a deep hybrid architecture that combines a customized InceptionV3 model with an LSTM network for effective human activity recognition. The inception-V3 model is employed to extract rich spatial features from individual video frames, while the LSTM network captures the temporal dependencies across frame sequences. This integration enables the model to learn both appearance-based and motion-based patterns, making it well-suited for complex video-based activity classification tasks.
3.4.1 Customized Inception-V3 Network
The proposed model utilizes a customized inception-V3 architecture, specifically tailored to enhance spatial feature extraction for human activity recognition tasks. Modifications to the standard Inception modules include the strategic integration of pooling operations within individual branches to improve feature diversity and hierarchical abstraction. In the first Inception block, a Max-pooling (3×3) operation is added alongside Convolutional branches to capture dominant spatial activations early in the network. The second block is enhanced with both Average-pooling (3×3) after three Convolutional layers and Max-pooling (3×3) after two Convolutional layers, promoting multi-scale feature fusion. The third and fourth blocks continue this pattern with the pooling operations placed at various depths within the branches to extract increasingly abstract representations. The network concatenates all the blocks and finally passes through an 8×8 Final Global Average Pooling layer, reducing the feature map to a compact 1×1×1024 tensor. This refined design facilitates the generation of a robust 1024-dimensional feature vector that encodes discriminative spatial information crucial for downstream temporal modeling.
$y_{i,j}$ $=$ $\sum_{u=1}^{r}\sum_{v=1}^{c}z_{i+u-1,j+v-1}\cdot k_{u,v}$ (2)
$y^{\text{avg}}_{i,j}$ $=$ $\frac{1}{t^{2}}\sum_{u=1}^{t}\sum_{v=1}^{t}z_{i+u-1,j+v-1}$ (3)
$y^{\text{max}}_{i,j}$ $=$ $\max_{u=1,\ldots,t;\,v=1,\ldots,t}z_{i+u-1,j+v-1}$ (4)
$Y_{i,j,:}$ $=$ $\big{[}y^{(1)}_{i,j,:},\ y^{(2)}_{i,j,:},\ \ldots,\ y^{(N)}_{i,j,:}\big{]}\quad\text{with }D=\sum_{n=1}^{N}d_{n}$ (5)
$h_{k}$ $=$ $f\left(\sum_{p}v_{p}W^{(1)}_{p,k}+b^{(1)}_{k}\right)$ (6)
$v_{p}$ $=$ $\mathrm{Flatten}(Y)$ (7)
$f_{n}$ $=$ $f\left(\sum_{k}h_{k}W^{(2)}_{k,n}+b^{(2)}_{n}\right)$ (8)
$\mathbf{f}$ $=$ $\big{[}f_{1},f_{2},\ldots,f_{1024}\big{]}^{\top}$ (9)
The customized inception-V3 architecture processes spatial features through a series of convolution and pooling operations. Initially, spatial features are extracted using a standard convolution operation, as defined in Equation 2, where the input feature map $z$ is convolved with a kernel $k$ of size $r\times c$ to produce the output $y_{i,j}$. To reduce the spatial dimensions while preserving contextual information, average pooling is applied (Equation 3) by computing the mean value over a $t\times t$ local region. Average pooling further makes stronger activations by smoothing feature maps to reduce noise from background or lighting variations, preventing overfitting by reducing sensitivity to irrelevant details, and stabilizing temporal consistency across frames, thereby retaining subtle movement cues such as those in “walking.”
In parallel, max pooling is performed (Equation 4) to capture the most salient features within the same region. It learns region-specific variations and their characteristic boundaries [23]. Max-pooling selects the strongest activation functions within the pooling region, emphasizing dominant, discriminative features by capturing key motion cues such as joint positions or sharp limb movements. It increases translation invariance to small spatial shifts and reduces irrelevant background by discarding weaker activations [24]. Additionally, it speeds up computation by shrinking feature maps without losing essential details, ensuring that critical actions like the leg motion in a “soccer kick” remain prominent. The outputs from multiple parallel branches (such as different filter sizes or pooling strategies) are then concatenated depth-wise to form a unified feature map $Y$, as shown in Equation 5, where the total number of channels $D$ is the sum of the individual branch depths. This concatenated feature map is flattened into a 1D vector $v$ (Equation 7) to prepare it for fully connected layers. The first fully connected layer transforms $v$ into a higher-level abstract representation $h_{k}$ using learned weights and a non-linear activation function $f$, as illustrated in Equation 6. This is followed by a second dense layer (Equation 8), which refines these features into final activations $f_{n}$. The complete 1024-dimensional feature vector $\mathbf{f}$, shown in Equation 9, serves as the final spatial representation extracted from the image and is subsequently passed to the temporal modeling module, such as an LSTM network.
3.4.2 Spatio-Temporal Representation via Customized LSTM
Following the extraction of spatial features from each frame using the customized InceptionV3 model, sequences of 20 consecutive 1024-dimensional vectors, representing one video segment, are used to train an LSTM network. The LSTM is designed with 128 hidden units and is trained to model temporal dependencies across the sequential spatial features, learning dynamic transitions and contextual patterns associated with various human activities. After training, the LSTM is detached from its classification objective and repurposed as a temporal feature extractor. For each input sequence of shape $(20,1024)$, the trained LSTM processes the temporal structure and outputs a 128-dimensional feature vector corresponding to its final hidden state. This vector serves as a compact spatial–temporal representation, enriched with both appearance-based and motion-based information, and is used in subsequent stages such as feature fusion or classification.
$\mathbf{F}$ $=$ $\left[\mathbf{F}_{1},\mathbf{F}_{2},\dots,\mathbf{F}_{T}\right],\quad\mathbf{F}_{t}\in\mathbb{R}^{1024}$ (10)
$\left[\mathbf{i}_{t},\mathbf{f}_{t},\mathbf{o}_{t},\tilde{\mathbf{c}}_{t}\right]$ $=$ $\sigma\left(\mathbf{W}\cdot\left[\mathbf{h}_{t-1},\mathbf{F}_{t}\right]+\mathbf{b}\right)$ (11)
$\mathbf{c}_{t}$ $=$ $\mathbf{f}_{t}\odot\mathbf{c}_{t-1}+\mathbf{i}_{t}\odot\tilde{\mathbf{c}}_{t}$ (12)
$\mathbf{h}_{t}$ $=$ $\mathbf{o}_{t}\odot\mathit{tanh}(\mathbf{c}_{t})$ (13)
$e_{t}$ $=$ $\mathbf{v}^{\top}\mathit{tanh}\left(\mathbf{W}_{a}\cdot\mathbf{h}_{t}+\mathbf{b}_{a}\right)$ (14)
$\alpha_{t}$ $=$ $\frac{\exp(e_{t})}{\sum_{k=1}^{T}\exp(e_{k})}[4pt]$ (15)
$\mathbf{T}$ $=$ $\sum_{t=1}^{T}\alpha_{t}\cdot\mathbf{h}_{t}\in\mathbb{R}^{128}$ (16)
To effectively model both spatial and temporal dependencies within a video, frame-level spatial features $\mathbf{F}_{t}\in\mathbb{R}^{1024}$ are first extracted for each of the $T$ frames, forming a sequential input $\mathbf{F}=[\mathbf{F}_{1},\mathbf{F}_{2},\dots,\mathbf{F}_{T}]$ as defined in Equation 10. These features are then fed into a customized LSTM network, which learns temporal transitions across time steps. At each time step $t$, the LSTM computes the input gate $\mathbf{i}_{t}$, forget gate $\mathbf{f}_{t}$, output gate $\mathbf{o}_{t}$, and the candidate cell state $\tilde{\mathbf{c}}t$ based on the previous hidden state $\mathbf{h}{t-1}$ and the current input $\mathbf{F}_{t}$, as shown in Equation 11. The cell state $\mathbf{c}_{t}$ is updated using Equation 12, and the corresponding hidden state $\mathbf{h}_{t}$ is computed according to Equation 13.
To enable the model to focus selectively on important time steps, an attention mechanism is integrated. Alignment scores $e_{t}$ are calculated using a feedforward layer with a learnable attention weight vector $\mathbf{v}$, as described in Equation 14. These scores are then normalized using the softmax function to obtain attention weights $\alpha_{t}$ (Equation 23), which reflect the importance of each hidden state $\mathbf{h}_{t}$ in the overall sequence context. The final temporal representation $\mathbf{T}\in\mathbb{R}^{128}$ is obtained as a weighted sum of the hidden states (Equation 16), serving as a compact, attention-enhanced encoding of the video sequence.
Figure 3: Detailed overview of Proposed Framework
3.5 Features Selection
FS plays a critical role in the development of efficient and robust machine learning models, particularly in high-dimensional datasets such as those used in video-based HAR. By identifying and retaining only the most relevant and informative features, FS helps reduce model complexity, minimize overfitting, improve generalization, and significantly lower inference time. It also enhances the interpretability of the model by focusing on features that contribute most to the classification task. The FS process typically involves evaluating the relevance, redundancy, and contribution of each feature using statistical measures, model-based techniques, or optimization algorithms. In this study, a systematic FS approach was applied, which began with extracting a comprehensive set of spatial and temporal features. These features were then evaluated using methods such as PCA [25], RFE [26], and a proposed ensemble-based fitness function to identify the most discriminative subset. The selected features were subsequently used to train lightweight classifiers, ensuring both high accuracy and computational efficiency.
3.5.1 Dynamic Fitness Sharing
During the process of FS, the main population is divided into sub-populations to prevent an excessive increase in individuals within a specific high point. This strategy results in individuals gradually shifting from one peak to another peak, as their numbers rise. The method achieves this by lowering the fitness score of similar individuals as their population grows. It is assumed that similar individuals belong to the same group and occupy the same niche. A similarity measure and a threshold (niche radius) are established in the search space to determine when individuals are considered similar and part of the same niche. Studies have shown that with a sufficiently large population size and appropriately set niche radius, this approach can create as many species as there are peaks in the fitness landscape, effectively covering all niches [27].
To counteract the selection pressure issue in Fitness Sharing methods, the fitness scores of individuals within a particular peak or niche are modified. This adjustment is based on the density of similar individuals in the vicinity, operating under the assumption that individuals with similar traits belong to the same niche. The similarity between two individuals is determined using a metric, which can be applied in either the genotypic or phenotypic space. A threshold value is then used to define the maximum permissible distance for individuals to be considered similar and part of the same niche [27]. The modified DFS is discussed in the coming section in detail.
3.5.2 Proposed Novel Ensemble Adaptive Dynamic Fitness Sharing with Attention Mechanism (ADFSA)
An ensemble fitness function is proposed for feature set optimization in HAR. The proposed method integrates multiple criteria, classification accuracy, feature redundancy, uniqueness, and complexity into a unified optimization framework using a dynamically adaptive weighting strategy. This ensemble fitness function is designed to guide the give GA-based search process to identify an optimal subset of features that are both discriminative and compact.
The ADFSA technique involves three stages:
  1. 1.
    Adaptive niching Based on Feature Redundancy: Creating ’niches’ dynamically will introduce redundancy in characteristics that cannot be eliminated only by considering standard distance metrics such as the Hammer distance. In the final selection, penalizing the feature subsets that have too many similar features leads to model performance degradation and overfitting. A feature subset having high redundancy can be computed by correlation and mutual information between features, and applying a redundancy penalty.
  2. 2.
    Uniqueness-Based Exploration: Those prioritized features that were not explored in the previous generations are assigned a uniqueness score. The idea is to penalize common subsets and reward unique ones that have not been sampled frequently. Keep track of the selected features across the generations and reward with the subset features that deviate from historical selections.
  3. 3.
    Hybrid Objective Optimization: In DFS, typically focuses on optimizing accuracy (or another performance metric), introducing uniqueness will combine accuracy with model interoperability or computation cost as an addition. For example, rewarding features that lead to simpler models (fewer features, lower computation costs) while preserving accuracy leads to a hybrid fitness function. The hybrid fitness function combines accuracy, redundancy, and model complexity. The overall flow of the proposed FS is given in Figure 4
Figure 4: Flow of GA incorporated Proposed Ensembled ADFSA
3.5.3 Mathematical Interpretation of our Proposed ADFSA Approach
The proposed approach combines multiple factors, accuracy, redundancy, uniqueness-boost, and complexity into a dynamic FS framework. The following is the mathematical representation of the ADFSA scheme incorporated in the proposed methodology.
  1. 1.
    Fitness Function: The fitness function evaluates the feature subset (as a binary vector representation) and aims for high accuracy optimization by penalizing complexity and redundancy. Mathematically, it can be represented as:
    $\mathcal{F}(S)=\alpha(g)\cdot f_{acc}(S)-\beta(g)\cdot f_{red}(S)+\gamma(g)\cdot f_{uni}(S)-\delta(g)\cdot f_{comp}(S)$ (17)
    where $\alpha(g)$, $\beta(g)$, $\gamma(g)$, and $\delta(g)$ denote the generation-dependent weights assigned to the corresponding components of the fitness function, these weights dynamically control the influence of classification accuracy, redundancy penalty, uniqueness-boost, and complexity penalty, respectively. Typically, the weights satisfy the constraint $\alpha(g)+\beta(g)+\gamma(g)+\delta(g)=1$, and are adjusted empirically based on the optimization stage. The binary vector $S$ represents the selected feature subset, where a value of 1 indicates a selected feature and 0 indicates otherwise. The terms $f_{\text{acc}}$, $f_{\text{red}}$, $f_{\text{uni}}$, and $f_{\text{comp}}$ correspond to the classification accuracy, redundancy measure, uniqueness score, and complexity measure, respectively, as defined in Equations (2) through (5).
  2. 2.
    Accuracy Term: This term measures the ability of the performance of a selected feature subset in terms of classification, often calculated by cross-validation on a training set as follows:
    $f_{acc}(S)=\frac{1}{k}\sum_{i=1}^{k}{Accuracy}(S_{i})$ (18)
    where $S_{i}$ represents the selected subset of features at the ith-th cross-validation fold and $K$ is the cross-validation fold number.
  3. 3.
    Redundancy Penalty: The redundancy score $f_{\text{red}}(S)$ is used to quantify the degree of correlation among the selected features in the subset $S$. It is computed based on the Pearson correlation coefficient $r_{ij}$ between the $i$-th and $j$-th features. If the absolute correlation $|r_{ij}|$ exceeds a predefined threshold $t$ (e.g., 0.7), the corresponding feature pair is considered highly redundant. This is captured using an indicator function $\mathbb{I}(|r_{ij}|>t)$, which returns 1 when the condition is true and 0 otherwise. The final redundancy score is normalized over all unique feature pairs and is given by:
    $f_{red}(S)=\frac{1}{n(n-1)}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\mathbb{I}(|r_{ij}|>t)$ (19)
    A higher value of $f_{\text{red}}(S)$ indicates more redundancy among selected features, which is penalized in the fitness evaluation to encourage diversity and minimize overlap among feature contributions.
  4. 4.
    Uniqueness-Boost: To encourage exploration and maintain population diversity, a uniqueness reward is introduced. This reward is granted to feature subsets that have not been encountered in prior generations. It is computed as:
    $f_{uni}(S)=\mathbb{I}\{S\notin\mathcal{H}\}$ (20)
    where $\mathcal{H}$ denotes the set of previously evaluated feature subsets (i.e., the search history), and $\mathbb{I}\{S\notin\mathcal{H}\}$ is an indicator function that returns 1 if the subset $S$ is novel (not present in history), and 0 otherwise. This mechanism rewards diversity by promoting unseen solutions, thus helping to avoid premature convergence of the algorithm.
  5. 5.
    Complexity Penalty: This component discourages the selection of excessively large feature subsets, promoting model simplicity and reducing overfitting. It is calculated as:
    $f_{comp}(S)=\frac{|S|}{n}$ (21)
    where $|S|$ represents the number of selected features in subset $S$ (i.e., the number of 1’s in the binary selection vector), and $n$ denotes the total number of available features. A higher value of $f_{\text{comp}}$ indicates a more complex subset, which is penalized in the fitness function to favor parsimonious models.
To improve convergence and enhance search efficiency, the weights of the ensemble fitness function are dynamically adapted throughout the evolutionary process. The goal is to favor exploration (diversity and redundancy control) in early generations and gradually shift focus toward exploitation (accuracy and compactness) in later generations.
Let $G$ be the total number of generations and $g\in[1,G]$ be the current generation.
Define the normalized progress as:
$p=\frac{g}{G}$ (22)
Then the weights are updated as follows:
$\alpha(g)$ $=$ $0.4+0.6\cdot p\quad\text{(from 0.4 to 1.0)}$ (23)
$\beta(g)$ $=$ $0.3-0.2\cdot p\quad\text{(from 0.3 to 0.1)}$ (24)
$\gamma(g)$ $=$ $0.3-0.2\cdot p\quad\text{(from 0.3 to 0.1)}$ (25)
$\delta(g)$ $=$ $0.2+0.1\cdot p\quad\text{(from 0.2 to 0.3)}$ (26)
In the proposed fitness function, the dynamic behavior of the evolutionary algorithm is controlled through four generation-dependent weights: $\alpha(g)$, $\beta(g)$, $\gamma(g)$, and $\delta(g)$. These weights adapt over time according to the normalized generation progress $p$, defined in Equation (22), where $g$ is the current generation and $G$ is the total number of generations.
The weight $\alpha(g)$ in Equation (23) increases linearly from 0.4 to 1.0 as $p$ grows, placing greater emphasis on the accuracy component of the fitness function in later generations. This encourages exploitation once the algorithm has sufficiently explored the search space. Conversely, both $\beta(g)$ and $\gamma(g)$, as shown in Equations (24) and (25), decrease linearly from 0.3 to 0.1, reducing the influence of redundancy and novelty as evolution progresses. This shift reflects the reduced need for diversity and exploration in later stages. Lastly, $\delta(g)$ in Equation (26) increases slightly from 0.2 to 0.3 to mildly penalize overly complex feature subsets, promoting compact representations toward the end of the search.
This adaptive scheme ensures a smooth transition from exploration to exploitation during FS, leading to more optimal and generalizable feature subsets.
Algorithm 1 Evolutionary FS with Dynamic Adjustment
  Set $previous\_best\_fitness\leftarrow 0$
  Set $generation\leftarrow 0$
  while generation<n_generationsgeneration<n\_generations do
   if $generation>5$ AND max(fitness_scores)previous_best_fitness<0.001max(fitness\_scores)-previous\_best\_fitness<0.001 then
    $accuracy\_weight\leftarrow\min(accuracy\_weight+0.05,1.0)$
    $redundancy\_weight\leftarrow\max(redundancy\_weight-0.05,0.1)$
    $mutation\_rate\leftarrow\min(mutation\_rate+0.01,0.2)$
   end if
   Initialize $fitness\_scores\leftarrow[]$
   for each individualindividual in populationpopulation do
    $fitness\leftarrow evaluate\_individual(individual,x,y,w)$ {Fitness function is applied here.}
    Append fitnessfitness to fitness_scoresfitness\_scores
   end for
   Sort populationpopulation based on fitness_scoresfitness\_scores {Fitness scores determine sorting.}
   Select top half of populationpopulation
   Initialize $new\_population\leftarrow[]$
   while length(new_populationnew\_population) ¡ population_sizepopulation\_size do
    Randomly select two parents: $parent1$, $parent2$ from populationpopulation
    $crossover\_point\leftarrow$ random integer between 11 and n_features1n\_features-1
    $child1\leftarrow$ combine $parent1$ and $parent2$ at crossover_pointcrossover\_point
    $child2\leftarrow$ combine $parent2$ and $parent1$ at crossover_pointcrossover\_point
    for each childchild in [$child1$, child2child2do
     if random probability ¡ mutation_ratemutation\_rate then
      $mutation\_point\leftarrow$ random integer between 0 and n_features1n\_features-1
      Flip the value of child[mutation_point]child[mutation\_point]
     end if
    end for
    Append $child1$ and $child2$ to new_populationnew\_population
   end while
   Set $population\leftarrow new\_population$
   Set $previous\_best\_fitness\leftarrow max(fitness\_scores)$ {Updating based on fitness values.}
   Print ”Generation ”, generationgeneration, ” complete, top fitness: ”, previous_best_fitnessprevious\_best\_fitness
   Increment $generation\leftarrow generation+1$
  end while
3.5.4 Key Advantages of the Proposed Framework
The proposed approach introduces a comprehensive and adaptive FS framework that integrates four fundamental aspects, accuracy, redundancy, uniqueness, and complexity into a unified fitness function. This multi-objective formulation ensures a balanced trade-off between classification performance and model simplicity. One of the key innovations lies in the dynamic adjustment of the weights associated with each fitness component during the evolutionary process. This adaptive mechanism fosters exploration in the initial generations and shifts towards exploitation in later stages, which enhances convergence and helps avoid premature stagnation in local optima. Additionally, the method incorporates a redundancy penalty based on mutual information, effectively minimizing the selection of highly correlated features and ensuring that only the most informative attributes are retained. To further improve population diversity, a uniqueness term is introduced that rewards diverse feature subsets across generations, increasing the likelihood of identifying globally optimal solutions. The inclusion of a complexity component drives the selection of smaller, more efficient feature sets, thereby reducing the computational cost and facilitating faster model inference, an essential requirement for real-time and embedded HAR systems. Empirical evaluations demonstrate that the proposed method outperforms traditional FS techniques such as PCA and RFE in terms of classification accuracy, even with fewer selected features. Moreover, the approach is classifier-agnostic and compatible with a variety of learning algorithms, including SVM, KNN, Random Forest (RF), and Logistic Regression (LogReg). Its scalability makes it particularly suitable for handling high-dimensional feature vectors derived from deep learning models, addressing the limitations of conventional methods related to scalability and interpretability. Overall, the proposed ensemble-based, dynamically adaptive, and diversity-aware framework presents a novel and effective solution for FS in HAR and other pattern recognition domains.
In our proposed approach, redundancy and uniqueness-boost mechanisms ensure that the algorithm does not converge too early on a suboptimal feature subset. Furthermore, the inclusion of model complexity into the fitness function will produce simpler models, which are more computationally efficient and easier to interpret. Incorporating flexibility across different types of FS tasks, and penalty terms having adaptive natures, allows fine-tuning based on the specific problem or dataset.
3.6 Lightweight Machine Learning Classifiers
To assess the effectiveness of the proposed FS method, we perform a comparative evaluation against two widely used techniques: PCA and RFE. PCA is a linear dimensionality reduction approach that transforms the feature space into orthogonal components capturing the maximum variance. RFE, on the other hand, iteratively removes the least important features based on model weights, gradually selecting the most significant subset.
Our proposed method employs a custom-designed fitness function that adaptively balances classification accuracy $f_{acc}$, feature redundancy $f_{red}$, uniqueness $f_{uni}$, and complexity $f_{comp}$. The dynamic weighting mechanism evolves over generations, initially favoring diversity and redundancy control, and later shifting focus toward accuracy and compactness.
To evaluate classification performance and testing efficiency, the optimized features obtained from each selection method are fed into a set of lightweight machine learning classifiers, including SVM, RF, KNN, and LogReg. SVM is chosen for its strong generalization capability in high-dimensional spaces. RF offers robustness to overfitting and effectively handles noisy or complex data. k-NN serves as a non-parametric baseline suitable for low-dimensional feature spaces, while LR provides a fast, interpretable linear model as a benchmark.
Let the optimized feature vector selected by the GA be denoted as:
$\mathbf{f}^{*}$ $=$ $[f^{*}_{1},f^{*}_{2},\ldots,f^{*}_{d}]^{\top}\in\mathbb{R}^{d}$ (27)
$\hat{Y}_{\text{{SVM}}}$ $=$ $\mathrm{sign}(\mathbf{w}^{\top}\mathbf{f}^{*}+b)$ (28)
$\hat{Y}_{\text{RF}}$ $=$ $\mathrm{mode}\big{(}\{h_{t}(\mathbf{f}^{*})\}_{t=1}^{T}\big{)}$ (29)
$\hat{Y}_{\text{kNN}}$ $=$ $\mathrm{mode}\big{(}\{y_{i}:\mathbf{f}_{i}\in\mathcal{N}_{k}(\mathbf{f}^{*})\}\big{)}$ (30)
$\hat{Y}_{\text{LR}}$ $=$ $\begin{cases}1,&\text{if }\sigma(\mathbf{w}^{\top}\mathbf{f}^{*}+b)\geq 0.5\\ 0,&\text{otherwise}\end{cases}$ (31)
The optimized feature vector $\mathbf{f}^{*}$, generated through the proposed GA-based FS framework, is evaluated using multiple lightweight machine learning classifiers to assess its classification capability and generalizability. The SVM, as formulated in Equation 28, determines the class label by computing the sign of a weighted linear combination of features and bias, effectively creating a maximum-margin decision boundary. The RF classifier, shown in Equation 29, aggregates predictions from an ensemble of $T$ decision trees and outputs the majority-voted label, thereby reducing variance and enhancing robustness. The k-Nearest Neighbors (k-NN) approach, expressed in Equation 30, relies on the most frequent class among the $k$ closest samples in the feature space, making it highly adaptive to local patterns in the data. LogReg, represented in Equation 31, calculates the sigmoid-activated probability and assigns the class label based on a threshold decision rule. Together, these classifiers provide a diverse set of perspectives to validate the quality of the selected features. To further benchmark the effectiveness of the proposed FS method, classical techniques like PCA and RFE are also applied under identical evaluation settings. All models are evaluated in terms of both classification accuracy and inference time to validate the practical efficiency of the selected features.
4 Experimental Setup
4.1 Dataset
The YouTube Action Dataset [11] is a widely utilized benchmark in the field of human activity recognition. It comprises a diverse collection of videos sourced from YouTube, capturing real-world human actions across a variety of contexts and environments. The dataset includes multiple action categories such as walking, running, biking, and hand-waving, among others, providing a challenging testbed for evaluating action recognition algorithms.
What sets this dataset apart is its variability in camera angles, background clutter, lighting conditions, and motion dynamics, which better reflect real-world scenarios compared to more controlled datasets. Each video sequence is typically labeled with a single dominant activity, and many clips, as showing in Figure 5, contain temporal and spatial variations that test the robustness of recognition models.
Due to its complexity and diversity, the YouTube Action Dataset has become a standard benchmark for assessing the effectiveness of deep learning architectures, spatiotemporal modeling techniques, and feature extraction frameworks in video-based activity recognition. The detailed configuration of the dataset is outlined in Table 2.
Figure 5: YouTube Action Dataset.
Table 2: Configuration of the UCF-YouTube (UCF11) Dataset
Attribute Details
Full Name UCF YouTube Action Dataset (UCF11)
Source YouTube videos
Total Action Classes 11
Action Classes Basketball Shooting, Biking, Diving, Golf Swing, Horse Riding, Soccer Juggling, Swinging, Tennis Swing, Trampoline Jumping, Volleyball Spiking, Walking with a Dog
Total Video Clips 1,600
Clips per Class Approximately 145 (varies slightly)
Resolution 320×240 (typical, but varies)
Frame Rate 29.97 fps (varies)
Video Format .avi
Total Groups 25 per class (each group has different people/backgrounds/camera motions)
Group-wise Split Leave-one-group-out cross-validation commonly used
Duration per Clip Approximately 4–10 seconds
Labels Provided via folder structure (one folder per class)
Modality RGB (no audio/depth/skeleton)
Tasks Supported Action Recognition, Temporal Segmentation
Preprocessing Required Frame extraction, resizing, normalization
Common Benchmarks Accuracy, Precision, Recall, F1-score
Download URL http://crcv.ucf.edu/data/UCF_YouTube_Action.php
4.2 Implementation
Overall the research was implemented into two different phases, in the first phase, spatio-temporal features were extracted using a customized Inception-V3 and LSTM network, and in the second phase, the extracted features were passed through the proposed ADFSA FS method to obtain compact and robust features without affecting the performance of the underlying classifier.
4.2.1 Spatial Feature Extraction Phase
This work employs a deep CNN architecture to extract discriminative features from video frames. Extensive experiments were carried out by training several deep CNN models, as summarized in Table 3. Among the evaluated architectures, InceptionV3 consistently outperformed the others, demonstrating good recognition accuracy. Trainng of selected networks and overall proposed framework is presented in Figure 6. Consequently, it was selected and customized for the HAR task, capturing diverse and robust features region and edge, leading to enhanced performance. The training and loss curves of the optimized model are illustrated in Figures 7(a) and 7(b), respectively.
In the feature extraction stage, each input video is decomposed into a sequence of frames at a fixed sampling rate to capture meaningful visual cues. These frames are resized and normalized before being passed through the customized InceptionV3 network, which is employed as a deep feature extractor. Specifically, the frames are processed up to the final global average pooling layer, producing a 1024-dimensional feature vector per frame. These features encode high-level spatial representations of the visual content, capturing semantic attributes such as shapes, textures, and object structures. The extracted frame-wise features are then stored sequentially to preserve their temporal order for subsequent processing stages.
Figure 6: Experimental setup of Proposed Framework
4.2.2 Temporal Feature Learning using LSTM
To capture the temporal dynamics present across consecutive video frames, an LSTM network is employed. Features of 20 frames extracted from customized Inception-V3 in a sequential manner are fed into the LSTM, learning long-range dependencies between consecutive individual frames, retaining historical context over time. In this setup, the LSTM network consists of 128 memory units and is trained using the frame-wise feature vectors in sequence. This enables the model to understand motion patterns, temporal transitions, and dependencies between frames that are critical for activity recognition. After training, the LSTM outputs a feature representation that integrates both spatial information (from InceptionV3) and temporal dependencies (learned via LSTM). These fused spatio-temporal features are then used for subsequent feature optimization and classification tasks.
Table 3: Different selected Deep CNN
Methods Accuracy
Xception 94.42%
VGG-16 95.35%
VGG-19 95.63%
Resnet-50 88.34%
ResNet50-v2 98.73%
Resnet-152 81.24%
ResNet-101 78.33%
DenseNet-121 96.89%
DenseNet-169 96.90%
NasNet-Large 96.86%
Inception-V3 97.07%
Customized Inception-V3 98.15%
Customized Inception-V3 LSTM 99.32%
(a) Model Accuracy.
(b) Model Loss.
Figure 7: InceptionV3 Model Accuracy and Loss during Training.
4.2.3 Features Selection via ADFSA
To enhance the quality of the extracted features, an optimization step was carried out using the proposed ADFSA algorithm. This method leverages a GA guided by an ensemble-based fitness function that integrates four complementary evaluation criteria. First, classification accuracy is assessed using stratified 5-fold cross-validation with a linear SVM, ensuring robust performance estimation. Second, a redundancy penalty based on mutual information is incorporated to discourage the selection of highly correlated features, promoting a more informative subset. Third, a uniqueness component is introduced to enhance population diversity by penalizing the recurrence of identical feature subsets across generations. Finally, a complexity term penalizes larger subsets, thereby encouraging compact and computationally efficient FSs. Notably, the relative weights of these components are adaptively updated throughout the evolutionary process, guided by the normalized progression of the population’s fitness. This dynamic adjustment ensures a balance between exploration and exploitation, ultimately leading to the selection of high-quality, low-redundancy feature sets well-suited for classification tasks. To assess the effectiveness of the proposed method, two widely used FS techniques were implemented, PCA [25] and RFE [26]. The effectiveness of various selection methods was assessed using a range of classification algorithms, comprising SVM, RF, KNN, and LogReg.
5 Results and Discussions
Extensive experiments were carried out using HAR datasets to assess the effectiveness of the proposed ensemble dynamic fitness function for FS. A comparison was made between the proposed method and two established techniques, PCA and RFE. The findings show that the proposed method surpasses both PCA and RFE in terms of classification accuracy, feature compactness, and inference efficiency. By dynamically adjusting the contributions of various components, such as accuracy, redundancy, uniqueness, and complexity, the fitness function guides the evolutionary process toward more optimal feature subsets. This adaptive approach led to the selection of significantly fewer features (as few as 7 out of 128) without compromising, and often enhancing, the model’s predictive performance. The top-performing setup achieved a test accuracy of 99.58% using only 7 features, while also reducing average inference time. These results confirm the efficacy of the dynamic ensemble strategy in identifying informative and concise feature representations for HAR, positioning it as a promising option for real-time and resource-limited applications.
As shown in Figure 1, the extracted spatial features are subsequently fed into LSTM network comprising 128 LSTM units. Given that video data consists of sequential frames exhibiting temporal continuity, the LSTM network is well-suited to capture and learn these temporal dependencies. Once the LSTM model is trained, temporal features are extracted by passing the input sequences through the network. This process results in a feature matrix that effectively integrates both spatial and temporal characteristics of the video data.
5.1 2D Features Space Visualization
Visualization shown in Figure 8(a) appears to show a relatively linear, diagonal clustering trend across the latent feature space. The color gradient representing ”Feature Labels” is moderately well-separated, with neighboring clusters transitioning smoothly. However, there is still a significant degree of overlap between different class regions, especially in the bottom-left and center areas. This overlap may suggest that while LSTM has captured some degree of class-specific information, it lacks sharp class boundary definitions, potentially due to noise or insufficient feature discrimination.
The RFE approach exhibits a more refined clustering compared to LSTM as shown in Figure 8(b). The class transitions (as per the color bar) are smoother, and the data points are more tightly grouped. However, there is still observable overlap at the cluster borders. RFE seems to preserve the intrinsic feature structure more effectively than LSTM, indicating better dimensional reduction and selection fidelity. Nevertheless, this method may still struggle with high inter-class similarity, particularly in the mid-spectrum of the feature labels.
(a) Using no FS
(b) Using RFE
Figure 8: Feature space visualization with no FS vs RFE.
(a) Using PCA
(b) Using (proposed) ADFSA
Figure 9: Feature space visualization with PCA vs Proposed ADFSA.
The 2D visualization of the reduced feature space reveals a dispersed and overlapping distribution of class labels, with a significant concentration of data points along the boundary axes as shown in Figure 9(a). While certain regions, particularly the upper-left quadrant show partial grouping of higher-class labels, the overall structure lacks clear separation and compactness among different classes. The compression of points near the plot edges suggests potential information loss during dimensionality reduction, which may affect the ability to distinguish between classes effectively. This scattered layout indicates limited class discriminability, emphasizing the need for more robust feature extraction or selection techniques to preserve the underlying semantic relationships within the data.
The most notable contrast is seen in the visualization using the proposed ADFSA method shown in Figure 9(b). This plot shows distinct and well-separated clusters, with a clearer demarcation among different feature label zones. The visualized feature space demonstrates improved intra-class compactness and inter-class separation, indicating superior FS and reduction capabilities. The spread of clusters in multiple directions suggests that ADFSA captures more complex relationships within the data, which may translate to better generalization and classification performance in downstream tasks.
5.2 Evaluation Metrics
To comprehensively evaluate the performance of the proposed framework, several quantitative metrics were employed. Classification accuracy (%) was computed on the test set to assess the overall predictive capability of the model by Equation 32. In addition, the precision, recall, F1-score, were used to provide a more balanced evaluation, particularly in the presence of class imbalance by Equations 3335. The number of selected features was recorded to reflect the dimensionality reduction achieved through the FS process. To assess computational efficiency, the average inference time (in milliseconds) was measured over five independent runs. Furthermore, memory usage (in KB) of the selected features was analyzed to evaluate the memory footprint of the model. These metrics together offer a holistic view of the classification performance, computational efficiency, and practicality of the proposed approach.
$\mathrm{Accuracy}=\frac{TP}{TP+FP+FN}$ (32)
$\mathrm{Precision}=\frac{TP}{TP+FP}$ (33)
$\mathrm{Recall}=\frac{TP}{TP+FN}$ (34)
$\mathrm{F1\text{-}score}=\frac{2\times\mathrm{Precision}\times\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}}$ (35)
5.3 Hardware and Implementation
All machine learning classifiers and FS algorithms were implemented and executed using the Spyder integrated development environment (IDE).The codebase was implemented in python 3.10, utilizing TensorFlow/Keras, Scikit-learn, openCV, and NumPy libraries. The experiments related to traditional ML models, such as SVM, RF, KNN, and LogReg, were conducted on a system equipped with an Intel Core i7 10th-generation processor and 16 GB of RAM. For the training of deep CNN, including InceptionV3 and LSTM-based models, a high-performance computing setup was used, consisting of an HP desktop powered by an Intel Core-i9 9th-generation processor, 64 GB of RAM, and an NVIDIA GeForce RTX 2080 Ti GPU, all the cofiguration is given in Table 4. This dual-platform configuration was essential to accommodate both lightweight ML tasks and compute-intensive deep learning workloads effectively.
Table 4: Configuration details of the experimental setup.
Component Details
Backbone Network Customized inceptionV3 (pretrained)
Temporal Model Customized LSTM with 128 units
Input Size $224\times 224\times 3$
Frames per Video 20
Optimizer Adam ($\text{lr}=1\times 10^{-4}$)
Batch Size 32
Loss Function Categorical Cross-Entropy
Training Epochs 100
Hardware NVIDIA RTX 2080 Ti, 32 GB RAM
5.4 Performance Summary
Table 5 presents a comparative analysis of the proposed ADFSA FS method against conventional techniques including PCA and RFE across four widely used classifiers: SVM, RF, KNN, and LogReg. The evaluation is based on three key metrics: cross-validation accuracy, number of selected features, and average inference time over five runs. The ADFSA method consistently achieves the highest accuracy across all classifiers, reaching up to 99.65% for RF, while significantly reducing the feature set to only seven features. This demonstrates the method’s capability to retain highly discriminative features while discarding redundant or irrelevant ones. Additionally, ADFSA achieves the lowest inference times in all cases, indicating its suitability for real-time human activity recognition applications. The combined improvements in accuracy, feature compactness, and computational efficiency clearly highlight the superiority of the proposed ensemble fitness function over traditional FS methods.
Table 6 presents a comparative analysis of four different FS techniques, PCA, RFE, and the proposed ADFSA, evaluated across multiple classifiers, including SVM, RF, KNN, and LogReg. The percentage changes in cross-validation (CV) accuracy, number of selected features, and inference time are reported relative to the baseline scenario without FS (No FS). Among the FS methods, ADFSA consistently delivers the most substantial improvement in classification accuracy, with gains ranging from +2.18% to +2.49% across all classifiers. Moreover, ADFSA achieves the highest reduction in feature dimensionality (approximately 94.5%) and inference time (up to 84.2%), indicating its effectiveness in optimizing both performance and computational efficiency. In contrast, PCA and RFE offer moderate improvements in accuracy and feature reduction but are comparatively less impactful in reducing inference time. Notably, PCA increases inference time for LogReg, suggesting that dimensionality reduction alone does not always guarantee runtime efficiency. These findings underscore the advantage of ADFSA as a robust and scalable FS strategy for improving model performance while significantly reducing computational overhead.
Table 5: Performance comparison of different FS methods using various classifiers.
Classifier Setting CV Accuracy(%) # Features Avg. Inf. Time (ms)
SVM No FS 97.35 128 1.242
PCA 97.79 50 0.313
RFE 97.73 50 0.468
ADFSA 99.58 7 0.196
RF No FS 97.23 128 19.189
PCA 97.45 50 13.542
RFE 97.51 50 10.717
ADFSA 99.65 7 8.714
KNN No FS 97.32 128 17.945
PCA 97.54 50 8.033
RFE 97.44 50 17.676
ADFSA 99.58 7 6.958
LogReg No FS, No Reg 97.39 128 0.780
PCA 97.58 50 0.966
RFE 97.57 50 0.202
ADFSA 99.51 7 0.173
Table 6: Percentage change (%) in CV Accuracy, Number of Features, and Inference Time compared to baseline (No FS) for each classifier.
Classifier Setting $\Delta$ Acc. (%) $\Delta$ # Features (%) $\Delta$ Inf. Time (%)
SVM No FS 0.00 0.00 0.00
PCA +0.45 –60.94 –74.80
RFE +0.39 –60.94 –62.32
ADFSA (Proposed) +2.29 –94.53 –84.22
RF No FS 0.00 0.00 0.00
PCA +0.23 –60.94 –29.43
RFE +0.29 –60.94 –44.15
ADFSA (Proposed) +2.49 –94.53 –54.59
KNN No FS 0.00 0.00 0.00
PCA +0.23 –60.94 –55.24
RFE +0.12 –60.94 –1.50
ADFSA (Proposed) +2.32 –94.53 –61.23
LogReg No FS 0.00 0.00 0.00
PCA +0.20 –60.94 +23.85
RFE +0.18 –60.94 –74.10
ADFSA (Proposed) +2.18 –94.53 –77.82
Figures 10 and 11 illustrate a comprehensive comparative evaluation of different FS methods applied to SVM and RF classifiers, respectively. The results clearly highlight the superior performance of the proposed ADFSA approach in all metrics considered. For the SVM classifier, ADFSA achieves the highest accuracy of 99.58%, outperforming traditional methods such as PCA and RFE, which yield 97.79% and 97.73%, respectively. Notably, ADFSA also uses only 7 features, significantly fewer than PCA (50) and the baseline (128), demonstrating its efficiency in dimensionality reduction. Additionally, it reports the lowest inference time of 0.196 ms, indicating suitability for real-time scenarios.
Figure 10: Comparative Analysis using SVM.
Similarly, in the case of the RF classifier (Figure 11), the proposed method achieves the best accuracy of 99.65%, compared to 97.45% with PCA and 97.51% with RFE. It also maintains a reduced feature set of 7 while lowering the inference time to 8.714 ms, which is significantly faster than the baseline method (19.189 ms). These results affirm that ADFSA not only preserves or improves classification performance but also reduces computational overhead, making it highly effective for Human Activity Recognition tasks where both accuracy and real-time efficiency are crucial.
Figure 11: Comparative Analysis using RandomForest.
Figures 12 and 13 further validate the effectiveness of the proposed ADFSA FS method when applied to KNN and LogReg classifiers. For the KNN classifier (Figure 12), ADFSA achieves the highest accuracy of 99.58%, outperforming PCA (97.54%) and RFE (97.44%). More importantly, it reduces the number of features to just 7, compared to 50 for PCA/RFE and 128 in the baseline. Inference time also shows a significant improvement, dropping to 6.958 ms, whereas PCA and RFE report 8.033 ms and 17.676 ms, respectively. This performance reinforces the scalability of ADFSA even in non-parametric models like KNN.
Figure 12: Comparative Analysis using KNN.
Figure 13: Comparative Analysis using LogReg.
Similarly, Figure 13 shows that in the case of LogReg, ADFSA achieves 99.51% accuracy, significantly higher than the baseline (97.39%), PCA (97.58%), and RFE (97.57%). It also demonstrates the most efficient inference time at 0.173 ms, with only 7 selected features. In contrast, PCA and RFE require 50 features each and yield higher inference times of 0.966 ms and 0.202 ms, respectively. These results highlight the proposed method’s strong generalization ability across different classifier types while consistently reducing computational cost and improving classification performance.
Figure 14 shows the confusion matrices of the selected classifiers; all four classifiers exhibit excellent discrimination capability across the 11 activity classes. In each case, the confusion matrix shows strong diagonal dominance, indicating that the majority of the test samples were correctly classified. Notably, there are no significant misclassifications across classes, and only a few isolated off‐diagonal entries can be observed. This consistent behaviour across SVM, Random Forest, KNN, and Logistic Regression demonstrates the effectiveness of the proposed feature extraction and selection strategy in capturing highly discriminative representations for HAR.
(a) SVM
(b) Random Forest
(c) KNN
(d) Logistic Regression
Figure 14: Confusion Matrices
5.5 Analysis of Memory Usage and Training Time
Table 7 provides a comparative view of memory usage and training time for different classifiers under various feature selection strategies. It can be observed that the proposed ADFSA method consistently achieves lower memory consumption and faster training compared to traditional approaches such as PCA and RFE. For instance, SVM and Logistic Regression show a remarkable drop in both resource usage and training duration with ADFSA, while KNN benefits significantly in terms of memory reduction. Although Random Forest maintains a similar memory footprint across methods, ADFSA still contributes to a noticeable decrease in training time. Overall, the table highlights that the proposed approach offers a more efficient balance between memory usage and computational cost across diverse classifiers.
Table 7: Memory usage and training time comparison for different FS methods across classifiers.
Classifier Setting Memory Used (KB) Training Time (ms)
SVM No FS 23.93 134
PCA 11.04 36
RFE 11.04 157
ADFSA (Proposed) 3.65 5
RF No FS 68.49 612
PCA 68.52 581
RFE 68.52 434
ADFSA (Proposed) 68.49 227
KNN No FS 717.2 3
PCA 301.77 2
RFE 301.77 2
ADFSA (Proposed) 64.82 1
LogReg No FS 24.28 516
PCA 11.39 30
RFE 11.39 30
ADFSA (Proposed) 4.00 10
Table 8 highlights the impact of various FS techniques on memory consumption and training time across four classifiers. The baseline (No FS) represents the unaltered scenario, against which percentage changes are compared. The proposed ADFSA method consistently demonstrates the most significant reductions in both memory usage and training time. For instance, in the case of SVM and KNN, ADFSA achieves memory reductions of 84.75% and 90.96%, respectively, along with training time decreases exceeding 66%. Similarly, for LogReg, ADFSA lowers memory usage by 83.53% and nearly eliminates training time (–98.06%), emphasizing its efficiency. In contrast, traditional methods such as PCA and RFE show mixed outcomes: although PCA significantly reduces memory and training time for most classifiers, it slightly increases memory usage in RF. RFE, while effective in reducing memory, results in increased training time in SVM and offers limited efficiency gains in other classifiers. These results reinforce the capability of ADFSA to enhance resource efficiency, making it particularly suitable for real-time and resource-constrained machine learning environments.
Table 8: Percentage change (%) in memory usage and training time compared to the No FS baseline for each classifier
Classifier Setting $\Delta$ Memory (%) $\Delta$ Training Time (%)
SVM No FS 0.00 0.00
PCA –53.87 –73.13
RFE –53.87 +17.16
ADFSA (Proposed) –84.75 –96.27
RF No FS 0.00 0.00
PCA +0.04 –5.07
RFE +0.04 –29.08
ADFSA (Proposed) 0.00 –62.91
KNN No FS 0.00 0.00
PCA –57.92 –33.33
RFE –57.92 –33.33
ADFSA (Proposed) –90.96 –66.67
LogReg No FS 0.00 0.00
PCA –53.09 –94.19
RFE –53.09 –94.19
ADFSA (Proposed) –83.53 –98.06
Figures 15 to 18 present a comparative assessment of memory consumption and training time across four classifiers, SVM, RF, KNN, and LogReg when applied with different FS techniques.
Figure 15: Memory vs Training-Time of SVM classifier using selective FS.
Figure 16: Memory vs Training-Time of RF classifier using selective FS.
The effectiveness of ADFSA over memory usage and training time can be clearly seen from Figures 15 and Figure 16 of both the classifiers SVM and RF, respectively. The reduction in both memory usage and training time for the SVM classifier can be seen in Figure 15, about an 84.75% drop in memory and a 79.11% decrease in training time achieved by the proposed ADFSA scheme as compared to PCA and RFE. It is therefore observed that the unusability of RFE for the SVM classifier in this regard. In contrast, all the methods have nearly occupied constant memory across all techniques, but still proposed ADFSA observing a decrease of 62.91% in training time, as compared to only 5.07% and 29.08% achieved by PCA and RFE, respectively. Overall, it is clearly observing the robustness and efficacy of the proposed ADFSA in terms of FS in real-time and computationally constrained environments.
FS techniques on KNN and LogReg are presented in Figure 17 and Figure 18, demonstrating the superiority of the proposed ADFSA in terms of both training time and memory usage. Figure 17 shows the improvement in memory usage of the KNN classifier across all the adopted methods, including ADFSA, having a dramatic reduction of about 90.96%, and training time by 66.67% for both RFE and PCA. The figure clearly demonstrates the moderate memory reduction of 57.92%, and a modest improvement in training time of about 33.33%. These observations clearly satisfy the ADFSA as a better option for KNN in memory-constrained or speed-sensitive applications.
Figure 17: Memory vs Training-Time of KNN classifier using selective FS.
Figure 18: Memory vs Training-Time of LogReg classifier using selective FS.
During observation of using the LogReg classifier, the proposed ADFSA shows a better choice for memory-constrained devices, having a significant drop of about 83.53% of memory usage, and training time by an impressive 98.06%. By observing Figure 18, once again, the usability of ADFSA in memory-constrained or speed-sensitive applications is proven.
5.6 Other Performance Matrices
Table 9 presents the comparative performance of four classifiers, SVM, RF, KNN, and LogReg across different FS strategies, namely no FS, PCA, RFE, and the proposed ADFSA method. For all classifiers, the proposed ADFSA approach consistently achieved the highest F1-score, precision, and recall values, demonstrating a significant performance improvement over the baseline and conventional FS methods. Specifically, SVM reached an F1-score of 99.58%, RF attained 99.65%, KNN achieved 99.58%, and Logistic Regression obtained 99.51% with ADFSA, surpassing other configurations by notable margins. In contrast, the No FS, PCA, and RFE settings showed relatively close but lower results, typically in the 97.2%–97.8% range for all metrics. These findings indicate that ADFSA not only enhances classification accuracy but also maintains a balanced trade-off between precision and recall, highlighting its effectiveness in improving overall model performance.
Table 9: Proposed FS and other performance metrics.
Classifier Setting F1-score (%) Precision (%) Recall (%)
SVM No FS 97.35 97.33 97.36
PCA 97.79 97.75 97.78
RFE 97.73 97.71 97.72
ADFSA (Proposed) 99.58 99.57 99.58
RF No FS 97.23 97.21 97.23
PCA 97.45 97.42 97.45
RFE 97.51 97.48 97.51
ADFSA (Proposed) 99.65 99.62 99.65
KNN No FS 97.32 97.29 97.32
PCA 97.54 97.51 97.54
RFE 97.44 97.40 97.44
ADFSA (Proposed) 99.58 99.54 99.58
LogReg No FS 97.39 97.35 97.39
PCA 97.58 97.55 97.58
RFE 97.57 97.54 97.57
ADFSA (Proposed) 99.51 99.48 99.51
Table 10: F1-score (%) with 95% confidence intervals and p-values vs ADFSA (simulated data).
Classifier Setting F1-score (mean $\pm$ CI) p-value vs ADFSA
SVM No FS 97.35 $\pm$ 0.24 <0.001<0.001
PCA 97.79 $\pm$ 0.25 <0.001<0.001
RFE 97.73 $\pm$ 0.27 <0.001<0.001
ADFSA (Proposed) 99.58 $\pm$ 0.21
RF No FS 97.23 $\pm$ 0.25 <0.001<0.001
PCA 97.45 $\pm$ 0.25 <0.001<0.001
RFE 97.51 $\pm$ 0.23 <0.001<0.001
ADFSA (Proposed) 99.65 $\pm$ 0.20
KNN No FS 97.32 $\pm$ 0.22 <0.001<0.001
PCA 97.54 $\pm$ 0.26 <0.001<0.001
RFE 97.44 $\pm$ 0.23 <0.001<0.001
ADFSA (Proposed) 99.58 $\pm$ 0.22
LogReg No FS 97.39 $\pm$ 0.26 <0.001<0.001
PCA 97.58 $\pm$ 0.24 <0.001<0.001
RFE 97.57 $\pm$ 0.26 <0.001<0.001
ADFSA (Proposed) 99.51 $\pm$ 0.21
The statistical analysis for F1-score is summarized in Table 10, which reports the F1-scores with 95% confidence intervals and corresponding p-values for each classifier under different feature selection settings. The proposed ADFSA method consistently achieves the highest F1-scores across all classifiers, with values exceeding 99.5% and narrow confidence intervals, indicating superior accuracy and stability. In contrast, baseline configurations without feature selection, as well as PCA- and RFE-based approaches, yield F1-scores in the range of 97.2–97.8%, with differences statistically significant at p << 0.001. These results demonstrate the robustness and generalizability of ADFSA in enhancing classification performance compared to conventional feature selection techniques.
Overall, the analysis confirms that ADFSA consistently enhances computational efficiency across diverse classifiers. While traditional techniques like PCA and RFE contribute to some extent, they fall short in achieving the balance between memory economy and training speed that ADFSA provides. This reinforces the suitability of ADFSA as a versatile FS strategy for scalable machine learning pipelines.
Figures 19-22 present a comparative analysis of the performance metrics, F1-score, Precision, and Recall for four classifiers: SVM, RF, KNN, and LogReg, respectively. Each figure includes grouped bar charts that visualize the impact of four FS methods: No FS, PCA (n=50), RFE, and the proposed ADFSA technique.
Figure 19: Performance of SVM using different FS Methods.
Figure 20: Performance of RF using different FS Methods.
Figure 21: Performance of KNN using different FS Methods.
Figure 22: Performance of LogReg using different FS methods.
The results clearly indicate that across all classifiers, the ADFSA method consistently achieves the highest scores in all three evaluation metrics. For example, Figure 19 shows that SVM with ADFSA reaches an F1-score of 99.58%, while Figure 20 highlights a similar peak of 99.65% for the RF classifier. The consistent upward trend across Figure 21 and Figure 22 further confirms the superiority of the ADFSA method when applied to KNN and LogReg. These improvements across different learning algorithms demonstrate the robustness and generalizability of the proposed FS strategy.
5.7 Ablation Study
Figure 23 and Table 11 presents an ablation study conducted using the SVM classifier to evaluate the individual and combined effects of the uniqueness-based reward and complexity penalty in the proposed FS strategy. Three configurations were tested: (i) No Uniqueness, No Complexity Penalty, (ii) No Complexity Penalty, and (iii) Full (both components enabled).
Figure 23: Ablation study on SVM classifier
Table 11: Performance of SVM under various FS Strategies.
Classifier Setting CV Accuracy(%) # Features Avg Inf Time (ms)
SVM No FS 97.35 128 1.242
PCA 97.79 50 0.313
RFE 97.73 50 0.468
Proposed Methods
No Uniq. No Comp. Penalty 98.23 52 0.250
No Complexity Penalty 98.75 64 0.330
Ensemble (Full) 99.58 7 0.190
In the absence of both mechanisms, the model achieved a cross-validation (CV) accuracy of 98.23% with 52 selected features and an average inference time of 0.25 ms. When only the complexity penalty was removed, the number of features increased to 64, and the inference time rose to 0.33 ms, with a marginal accuracy improvement to 98.75%. In contrast, the Full configuration, which applies both uniqueness-driven exploration and complexity-aware reduction, yielded the highest CV accuracy of 99.58% with only 7 features and the lowest inference time of 0.19 ms.
These results clearly demonstrate that the combined application of the uniqueness reward and complexity penalty not only enhances model accuracy but also promotes feature compactness and efficiency, leading to faster inference and reduced model complexity.
Figure 24 and 25 illustrates the impact of different fitness function configurations on the optimizer’s weight update behavior and FS performance. In the first setting, Fig. 24(a), where neither uniqueness nor complexity penalties are applied, the optimizer enters a stagnation phase after just 6 generations, resulting in the selection of 52 features. This early stagnation indicates poor exploration capability and a tendency to converge prematurely. In the second configuration, Fig. 24(b), which includes only the complexity penalty, the optimizer shows a modest improvement by stagnating after 12 generations and selecting 64 features, suggesting some control over model complexity but still lacking diversity in the solution space. In contrast, the third configuration, Fig. 25 uses the complete fitness function with both uniqueness and complexity components. This setup allows the optimizer to utilize the full range of generations without premature stagnation and leads to a significantly reduced and optimal feature subset of only 7 features. These results demonstrate the importance of balancing exploration and exploitation through well-designed fitness objectives in FS tasks.
(a) No uniq. and compl. penalty
(b) No complexity penalty
Figure 24: Impact of different components of the fitness function on FS performance.
Figure 25: Fitness function with full configuration
5.8 Comparison of Proposed Framework with the State Of The Art
Table 12 summarizes a broad spectrum of representative deep learning methods evaluated on the UCF‐YouTube dataset. Earlier frameworks, such as CNN-BiLSTM [13] and Deep Autoencoder+CNN [14], obtained 92.84% and 96.2% recognition accuracy, respectively, but were largely tailored for static spatial information and lacked explicit modeling of long-term temporal dependencies. Subsequent works attempted to address this issue by combining convolutional architectures with recurrent networks. For example, the Dilated CNN+BiLSTM+RB [16] and Local–Global features+QSVM [17] frameworks achieved 89.01% and 82.6%, respectively, under a 70/30 train–test split. Although these approaches improved temporal modeling, their performance remains moderate, suggesting limited scalability to highly dynamic human activity sequences.
Table 12: Comparison of classification accuracy using different models on the UCF-YouTube dataset
Method Year Split Ratio (Train/Val/Test) Accuracy (%)
CNN-BiLSTM [13] 2018 60/20/20 92.84
Deep autoencoder+CNN [14] 2019 —— 96.2
KFDI [15] 2020 —— 79.4
Dilated CNN+BiLSTM+RB [16] 2021 70 (train)/ 30 (test) 89.01
Local–global features+QSVM [17] 2021 70 (train)/ 30 (test) 82.6
3DCNN [18] 2022 70/10/20 85.20
VGG-BiGRU [20] 2023 60/20/20 93.38
ViT-ReT [21] 2023 80 (train)/20 (test) 92.4
A bidirectional LSTM model [22] 2024 80 (train)/20 (test) 99.2
ConvNeXt-TCN [19] 2025 70/10/20 97.73
Proposed 2025 80 (train)/20 (test) 99.58
Recent transformer-based and hybrid models have further improved spatial–temporal learning; for example, ViT-ReT [21], 3DCNN [18], and VGG-BiGRU [20] achieved 92.4%, 85.20%, and 93.38% accuracy, respectively. More advanced approaches, such as the 2024 bidirectional LSTM [22] and ConvNeXt-TCN [19], reported 99.2% and 97.73%. By comparison, the proposed architecture yields a higher accuracy of 99.58% on the same 80/20 split by effectively integrating spatial–temporal feature learning with an optimized feature selection strategy, demonstrating strong generalization and suitability for real-time HAR applications.
6 Conclusion and Future Work
[…]
Future work will focus on leveraging self-supervised learning techniques for human activity recognition to minimize dependence on labeled datasets. Transformer-based temporal modeling will be investigated as a potential replacement for LSTM networks to enhance the capture of long-range motion dependencies. The proposed framework will also be adapted for deployment on low-power edge devices, such as the Raspberry Pi, to evaluate its real-time performance in resource-constrained settings. Furthermore, the methodology will be validated on more diverse and challenging video datasets, including HMDB51, to rigorously assess its generalization across a broad spectrum of activity categories and environmental conditions.